OpenStudy (annie96):

Which ordered pair makes the equation true? 4x – 9y = –5 (Points : 1) (–5, –2) (–1, 1) (10, 5) (20, 4)

OpenStudy (anonymous):

Okay, let's plug them in shall we? 4(-5)-9(-2) = -5. Now does it or does it not equal -5?

OpenStudy (anonymous):

So you basically multiply like so: \[(4\times-5)-(9\times-2)= -5\]

OpenStudy (anonymous):

@Annie96 Are you there?

OpenStudy (annie96):

yes i'm here.

OpenStudy (anonymous):

Are you able to tell me if that multiplication problem is true or false?

OpenStudy (annie96):

true.

OpenStudy (anonymous):

Well, let's see. \[4\times-5=-20\] So -20 for the first part, now let's do the second part. \[9\times-2=-18\] So -18 for the second. Now let's do -20 - -18 Two negatives, so we turn this into positive. -20 + 18 = -2 Thus, it's false. Not true. You see now?

OpenStudy (annie96):

yes

OpenStudy (anonymous):

Are you able to input the (-1,1) like we just did, for x and y variables? The first number, replace the x with -1. Then the second, replace the y with 1. Then, do what we just did.

OpenStudy (annie96):

Yes the true answer is (10,5) 4(10)-9(5)=-5

OpenStudy (anonymous):

\[\Large correct, good job :)!!\]

OpenStudy (annie96):

Thank You. for you're help.

OpenStudy (anonymous):

You're welcome. To correctly close a question, choose the best response, then close the question. Thanks :).